OpenStudy (anonymous):

HHHHEEELLP ASAPPP!! PLZ I GIVe medal! –4 • (–6 + 5.7) = –4 • (–6) +( )• 5.7 fill in the blank. the parenthesis is the blank. plz help

OpenStudy (anonymous):

@Jack1

OpenStudy (anonymous):

@grex10

OpenStudy (anonymous):

@Moo_Moo17

OpenStudy (anonymous):

@UnkleRhaukus please help !!

OpenStudy (unklerhaukus):

\[a\cdot(b+c)=a\cdot b+a\cdot c\] \[-4 \cdot (-6 + 5.7) = -4 \cdot (-6) +(? ) \cdot 5.7\]

OpenStudy (unklerhaukus):

\[\large\color{brown}a\cdot(\color{seagreen}b+\color{cornflowerblue}c)=\color{brown}a\cdot \color{seagreen}b+\color{brown}a\cdot \color{cornflowerblue}c\\[3ex] \large \color{brown}{-4} \cdot (\color{seagreen}{-6} + \color{cornflowerblue}{5.7}) = \color{brown}{-4} \cdot (\color{seagreen}{-6}) +(\color{brown}? ) \cdot \color{cornflowerblue}{5.7}\]

OpenStudy (unklerhaukus):

@officialkeily, what is the ( ? ) ?

OpenStudy (unklerhaukus):

@officialkeily!

OpenStudy (anonymous):

i got it already its -4

OpenStudy (unklerhaukus):

good work

OpenStudy (anonymous):

can you help with one more please @UnkleRhaukus

OpenStudy (unklerhaukus):

ok, sure

OpenStudy (anonymous):

A restaurant keeps track of sales each week. According to the scatter plot, which amount is an estimate of sales that will be made in week 8? A. $1500 B. $1000 C. $2000 D. $500

OpenStudy (unklerhaukus):

Which of those options would you most likely expect as the hight of the next green point on the plot

OpenStudy (anonymous):

UM A

OpenStudy (unklerhaukus):

Yes, that looks like the best of those options

OpenStudy (anonymous):

ok thx